OpenStudy (anonymous):

State the period, amplitude, max/min values, range, domain, horizontal phase shift and vertical displacement. y= 2cos(-1/2x +90^@)-3

OpenStudy (tkhunny):

You should get to it. Where do you plan to start?

OpenStudy (anonymous):

simplifying it first

OpenStudy (anonymous):

can you do it for me?

OpenStudy (tkhunny):

I don't think so. The point of this exercise is that it is as complicated as possible. You may wish to factor the stuff inside the argument of the cosine function.

OpenStudy (anonymous):

can you just put this equation in the form of y = a sin[b(x - c)] + d

OpenStudy (anonymous):

stop making it so complicated

OpenStudy (anonymous):

are you there?

OpenStudy (tkhunny):

Observation #1 - You are asking for help. Observation #2 - The fact that you are asking for help suggests you need help. Rule #1 - The Code of Conduct prevents just doing it for you. I can't be in your exam! Rule #2 - A volunteer trying to help you gets to pick HOW to help you. I already told you to factor the stuff in the argument of the cosine. Do that. This is not only to solve the problem, but also to prove that you have sufficient algebra skills so that you can be expected to be able to solve the problem. You have \(-\dfrac{1}{2}x + 90º\). Factor out -1/2 and the entire expression will be in the desired form.

OpenStudy (tkhunny):

Then you are saying that you do not know how to factor out the -1/2? That's all it takes. You can do it. I am nearly certain.

OpenStudy (anonymous):

ok I factored it

OpenStudy (tkhunny):

Show the factoring, please.

OpenStudy (anonymous):

at least tell me if its right ?

OpenStudy (tkhunny):

Show it and I will.

OpenStudy (anonymous):

y= 2cos(-1/2(x-180^@))-3

OpenStudy (tkhunny):

That's pretty good. Excellent work. Okay, here we go. y = cos(x) Period: 2pi Amplitude: 1 Range: [-1,1] <== Min and Max Domain: All Real Numbers Phase Shift: None Vertical Shift: None Do you believe all that?

OpenStudy (anonymous):

ok Thank you

OpenStudy (tkhunny):

That's just the beginning. The unaltered cosine function. We have not yet approached your problem statement.

OpenStudy (anonymous):

do you want to help me with a new question ?

OpenStudy (anonymous):

are you there?

OpenStudy (tkhunny):

Haven't done the old one, yet.

OpenStudy (anonymous):

its fine thats all I need

OpenStudy (tkhunny):

Well, then, excellent! Since you did it yourself.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok this is the next one that I need help with: The graph of y = -3sin( 2x) is shifted to the left by 30 units and up 5 units. Write the new equation.

OpenStudy (tkhunny):

Up is easy. Add "+5" on the end.

OpenStudy (tkhunny):

left and right are a little trickier. Shall we use +30 or -30?

OpenStudy (anonymous):

so it would be y= 3sin(2x) +5 ?

OpenStudy (anonymous):

-30?

OpenStudy (tkhunny):

No, sorry, that will move it right. It's a little counter-intuitive. We just need to play with the argument. We have (2x) It would be a mistake to write (2x + 30) You tell me why.

OpenStudy (anonymous):

it would give you the horizontal translation instead

OpenStudy (tkhunny):

No, we what a horizontal translation of 30. (2x + 30) is a horizontal translation of only 15. Remember that factoring? (2x+30) = 2(x+15) To achieve a left shift of 30, we must use [2(x+30)]

OpenStudy (anonymous):

+5 then?

OpenStudy (tkhunny):

We already did that. Out on the end that is a vertical shift of 5. Inside the argument cos(2x) vs. cos(2(x+30)) is a left shift of 30

OpenStudy (anonymous):

ok what is the full equation ?

OpenStudy (tkhunny):

You have the first half. y= 3sin(2x) +5 Add the other piece.

OpenStudy (anonymous):

[2(x+30)] ?

OpenStudy (tkhunny):

That's the piece. y= 3sin(2(x+30)) +5 Done.

OpenStudy (anonymous):

thank you !

OpenStudy (tkhunny):

No worries. Glad we could get along.

OpenStudy (anonymous):

I have another one

OpenStudy (anonymous):

Describe how the graph of y+3 = 2sin (3x+90^@) compares to the graph of y = sin(x)

OpenStudy (tkhunny):

Two things to prepare: 1) Move that 3 off the y. It needs to be on the other side to be in standard form. 2) Yet another time, factor that argument.

OpenStudy (anonymous):

ok so the 3 would become -3

OpenStudy (tkhunny):

Right. And the factoring?

OpenStudy (anonymous):

ok I got it !

OpenStudy (anonymous):

My answer: y= 2sin(3(x+30^@))-3 ??

OpenStudy (tkhunny):

BTW to get the degree symbol, hold down your [alt] key, and while holding it down, type on your numeric keypad 0 1 8 6.

OpenStudy (tkhunny):

Not quite an answer, but it is now in standard form.

OpenStudy (anonymous):

ok thanks so its not right?

OpenStudy (tkhunny):

If you want it in standard form, you are done. If you wish to describe how it is different from sin(x), you have only just begun.

OpenStudy (anonymous):

No I need to describe it to sin(x) how do I do that?

OpenStudy (tkhunny):

Then do that. What are the four pieces? y= 2sin(3(x+30^@))-3 ?? '2' in front does what to the amplitude? "-3" in the back does what to the vertical? "3" in the argument does what to the period? "+ 30º" in the argument does what to the phase shift?

OpenStudy (anonymous):

but thats the equation I had

OpenStudy (tkhunny):

There is no objection to the equation. What do all the parts do? This is the question. gtg Good luck.

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

what are the functions of sin(x)